OpenStudy (anonymous):

Please Show all Work. Type the missing number. Hint: Use the zero property of multiplication to find the answer. 1,239 × 7,498 ×? = 0

OpenStudy (anonymous):

goodluck

OpenStudy (king):

is it this please check \[1239\times7498x=0\]

OpenStudy (king):

if it is that then since 1239 and 7498 are constants they will not change,so x is the only variable and since their product is 0 then x needs to be 0.....!!!!!!!!!!!1

OpenStudy (anonymous):

thanks

OpenStudy (king):

ure welcome!!!!!!!

OpenStudy (anonymous):

can you help me with this one? Type the missing number. Hint: Use the identity property of multiplication to find the answer. 9,086 × 8,523 ×? = 77,439,978

OpenStudy (anonymous):

9,086 × 8,523 × a = 77439978 therefore a = 77439978/(9086*8523)

OpenStudy (anonymous):

thanks woo216

OpenStudy (anonymous):

this is simple algebra make sure u learn it well

OpenStudy (anonymous):

x = 0.